When a smaller number is subtracted from a larger number the difference is 34. If the larger number is 30 less than 3 times the smaller number, what are the two numbers.
Answer by ikdeep(226)

Let the smaller number = s
Let the largest number = l
According to first condition i.e When a smaller number is subtracted from a larger number the difference is 34,,,,we get the equation
l - s = 34...............(1)

According to second condition i.e larger number is 30 less than 3 times the smaller number,,,,we get the equation
l = 3s - 30 ...or....
l - 3s = -30..............(2)
By applying elimination method in (1)(2) you can get the answer.

Still if you won't be able to solve further, contact me again.
hope this will help you
Please feel free to revert back for any further queries